OpenStudy (anonymous):

initial investment 3000 doubles value 8.3 years. Assuming compound continuously. What is the interest rate? Pe^rt

OpenStudy (kropot72):

We can write the following equation from the given information: \[\large 6000=3000e^{8.3t}\ ............(1)\] Dividing both sides of (1) by 3000 we get: \[\large 2=e^{8.3t}\ ............(2)\] Now we need to solve equation (2) to find the interest rate as a decimal. The first step in doing that is to take natural logs (ln) of both sides of (2). Can you do that?

OpenStudy (kropot72):

@sharonholmes1 Are you there?

OpenStudy (anonymous):

I here and that's where I got stuck.

OpenStudy (kropot72):

The left hand side is simple. For the right hand side you use the following rule of logs: \[\large \ln e^{x}=x\] So using that rule what is: \[\large \ln e^{8.3t}=?\]

OpenStudy (kropot72):

@sharonholmes1 Are you there?

OpenStudy (anonymous):

I'm here.

OpenStudy (kropot72):

Can you answer my last post please.

OpenStudy (anonymous):

I got 8.3%

OpenStudy (kropot72):

That is not the correct result. Taking natural logs of both sides of (2) we get: \[\large \ln 2=8.3r\ .........(3)\] from which the value of r is: \[\large r=\frac{\ln 2}{8.3}=you\ can\ calculate\] The value of r must be multiplied by 100 to become a percentage rate. Note: In my previous postings my equations stated t where it should have been r.

OpenStudy (anonymous):

So will it be 8.4% ?

OpenStudy (kropot72):

My result to 2 decimal places is 8.35%

OpenStudy (anonymous):

That's what I got the first time and you said it was not right. So I thought I had to round it or something...lol.

OpenStudy (kropot72):

Your post stated: "I got 8.3%" For interest rates taken continuously over a long period, the percentage rate would normally be shown to two decimal places. So if the rate was truly exactly 8.3% it should be shown as 8.30%.

OpenStudy (anonymous):

Sorry about that. One of the choices was just 8.3 so I didn't think that the decimal place would have matter.

OpenStudy (kropot72):

If the answer was required to one decimal place, then 8.35% would round to 8.4%.

OpenStudy (kropot72):

Is 8.4% an answer choice?

OpenStudy (anonymous):

Yes. I rounded it to the nearest tenth of a percent.